find the equation of the line that passes through the intersection of the given pair of lines and satisfies the other given condition. graph.
3x+4y-2=0, 3x-4y+1=0; the intercepts are equal

Answer by Edwin McCravy(20086)

find the equation of the line that passes through the intersection 
of the given pair of lines and satisfies the other given condition. 
graph. 

3x + 4y - 2 = 0, 3x - 4y + 1 = 0; the intercepts are equal.

First we must find the point of intersection by 
the system of two equations in two unknowns
 
3x + 4y - 2 = 0 
3x - 4y + 1 = 0

You can solve that system of equations and get
 x = 1/6, y = 3/8

Now the problem become:

Find the equation of the line through the point
(1/6, 3/8)
which has equal intercepts.  To have equal 
intercept, if the point ehere the line crosses 
the y-axis is (0, b), then the point where the 
line crosses the x-axis is (b, 0).

So we find its slope by the slope formula

(x1, y1) = (0, b) and (x2, y2) = (b, 0)

     y2 - y1     (0) - (b)     -b
m = --------- = ----------- = ---- = -1 
     x2 - x1     (b) - (0)      b

Now we use the point slope form with m = -1 
and (x1, y1) = (1/6, 3/8)

y - y1 = m(x - x1) 
 
y - (3/8) = -1(x - 1/6)

y - 3/8 = -x + 1/6

Now we multiply through by LCD = 24 to clear 
of fractions: 

24y - 9 = -24x + 4

24y + 24x - 13 = 0

or solve for y and get

y = -x + 13/24 

The two equal intercepts are at points 

(13/24, 0) and (0, 13/24)

Edwin